OpenStudy (qqstory):

Trig Substitution, I dont understand anti derivative of (sqrt(x^2-4x+1)).

OpenStudy (qqstory):

I dont understand how come the right side of the answer is -3/2 ln | (x-2)+sqrt(x^2-4+1) |+C

OpenStudy (qqstory):

http://puu.sh/fIu6p/28469dcffb.png

OpenStudy (qqstory):

This is the all the steps to the solution, but what happened to the sqrt(3) at the bottom?

OpenStudy (thomas5267):

It gets absorbed into C. \[ \ln\left(\frac{a}{b}\right)=\ln(a)-\ln(b) \]

OpenStudy (qqstory):

How does that work?

OpenStudy (qqstory):

is it because ln(b) is just some sort of constant that we don't care?

OpenStudy (qqstory):

In other words, are all constants absorb into C?

ganeshie8 (ganeshie8):

\(\ln \sqrt{3}\) is a constant

ganeshie8 (ganeshie8):

and you must be knowing that \(C\) is an "arbitrary" constant

ganeshie8 (ganeshie8):

\(C\) is as much arbitrary as \(\ln\sqrt{3} + C\) if you like, you may define a new constant : \(C' =\ln\sqrt{3} + C \) and use it

ganeshie8 (ganeshie8):

thats called lumping of constants

OpenStudy (qqstory):

I understand this now, thanks a lot thomas & ganeshie

ganeshie8 (ganeshie8):

here is a fun 0=1 fake proof using this lumping trick http://math.stackexchange.com/questions/806254/using-integration-by-parts-results-in-0-1